find the difference 4/10-1/5
find the sum 1/8+1/10

Answer by stanbon(75887)

find the difference 4/10-1/5
=(20-10)/50=10/50=1/5
-------------------
find the sum 1/8+1/10
=18/80 = 9/40
-------------------
Cheers,
Stan H.